A car starts from rest and accelerates at uniform rate of 6 metres per second square for some time then moves at constant speed for sometime and retards is the same winform rate and the total time for the journey is 24 seconds and average speed for journey is 20 metres per second how long does the car move with constant speed

Grade:11

1 Answers

Vikas TU
7 years ago
x1 = 0 + 0.5*6t1^2......................(1)
x2 = (6*t1)*t2..............(2) {at constant speed}
x3 = (6*t1)t3 – 0.5*6*t3^2...............(3) {During retardation}
 
t1 + t2 + t3 = 24...............(4)
avg. speed = total distance / total time
=> (x1 + x2 + x3)/24 = 20 ….....................(5)
solve all the eqns. and get each values.